### v2

# Joint Beamforming Design and Bit Allocation in Massive MIMO with Resolution-Adaptive ADCs

Abstract
Low-resolution analog-to-digital converters (ADCs) have emerged as a promising technology for reducing power consumption and complexity in massive multiple-input multiple-output (MIMO) systems while maintaining satisfactory spectral and energy efficiencies (SE/EE). In this work, we first present the fundamental properties of optimal quantization and leverage them to derive a more accurate approximation of the covariance matrix of the quantization distortion. This theoretical finding facilitates the analysis of the system’s SE in the presence of low-resolution ADCs. Then, considering resolution-adaptive ADCs, we focus on the joint optimization of the transmit-receive beamforming and bit allocation to maximize the SE under constraints on the transmit power and the total number of active ADC bits. To solve the resulting mixed-integer problem, we first develop an efficient beamforming design for fixed ADC resolutions. Subsequently, we propose a low-complexity heuristic algorithm to iteratively optimize the ADC resolutions and beamforming matrices. Numerical results for a $64\times 64$ MIMO system demonstrate that the proposed design offers $6\%$ improvements in both SE and EE with $40\%$ fewer active ADC bits compared with uniform bit allocation. Furthermore, it is unveiled that receiving more data streams with low-resolution ADCs can lead to higher SE and EE compared with receiving fewer data streams with high-resolution ADCs.
Index Terms:
Beamforming, bit allocation , energy efficiency, massive MIMO, low-resolution ADCs, spectral efficiency.
I Introduction
Massive multiple-input multiple-output (MIMO) is a crucial physical-layer technology for wireless communications at both sub-6GHz and millimeter-wave (mmWave) frequencies [heath2016overview], addressing the increasing demand for high data rates [jiang2021road]. The large number of antenna elements in massive MIMO provides significant spatial multiplexing gains through beamforming techniques. Digital beamforming (DBF) architectures, which deploy a dedicated radio-frequency (RF) chain for each antenna element, can enable high spectral efficiency (SE) but incur substantial energy costs due to power-intensive RF components, especially analog-to-digital converters (ADCs). For instance, a high-speed ADC operating at $1$ Gsample/s with high resolution (e.g., 888812121212 bits) can consume several Watts [li2017channel]. Furthermore, its power consumption increases linearly with the signal bandwidth and exponentially with the number of resolution bits [murmann2015race], posing a significant challenge to the system’s energy efficiency (EE). Consequently, the integration of low-resolution ADCs and DBF has emerged as an effective strategy to curtail power consumption without unduly compromising the SE [liu2019low].
Another attractive solution in this regard is to utilize hybrid beamforming (HBF) architectures, where a small number of RF chains is connected to the antenna array through a network of phase shifters or switches [mendez2016hybrid]. However, HBF architectures have limited multiplexing capabilities and strongly depend on the calibration of the analog components [roth2018comparison]. Consequently, DBF requires lower circuit cost to achieve a SE similar to that of HBF [yan2019performance], which makes the former more energy efficient, especially when using low-resolution ADCs [roth2018comparison, castaneda2021resolution]. Moreover, while the water-filling (WF) power allocation achieves the capacity of a full-resolution MIMO system with perfect channel state information (CSI) at both the transmitter and receiver [tse2005fundamentals], it becomes suboptimal in the presence of low-resolution ADCs, necessitating a more efficient design.
I-A Prior Works
Recent years have witnessed a proliferation of studies on massive MIMO with low-resolution ADCs (with 22224444 bits). It has been shown in [singh2009limits, jacobsson2017throughput] that a system using very few bits can approach the performance of a full-resolution one. Mezghani et al. [mezghani2012capacity] derived a closed-form lower bound for the capacity of a point-to-point MIMO system. More recent works have focused on beamforming designs [mezghani2009transmit, jacobsson2017quantized, ling2019performance, ma2024digital]. Furthermore, mixed-ADC systems, which simultaneously deploy one-bit and high-resolution ADCs, have been shown to perform better than fixed-resolution architectures, especially at high SNR [zhang2016mixed, zhang2017performance, pirzadeh2018spectral]. On the other hand, resolution-adaptive ADCs have been studied in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, prasad2020optimizing, castaneda2021resolution, castaneda2021spawc, sheng2020energy] to flexibly balance the SE-EE tradeoff of low-resolution systems. For instance, it has been shown in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, sheng2020energy] that efficient bit allocation strategies can offer a higher EE compared with uniform-resolution architectures. Moreover, Castañeda et al. [castaneda2021resolution] developed a resolution-adaptive fully digital receiver within an application-specific integrated circuit (ASIC). Furthermore, they demonstrated that a $256$-antenna base station with resolution-adaptive ADCs serving $16$ users allows a reduction in power consumption by $6.7$ times compared with a traditional fixed-resolution design [castaneda2021spawc]. Additional SE gains can be achieved by jointly optimizing the transmit power and ADC resolutions [prasad2020optimizing].
Among the aforementioned works, two primary methods are used to model quantization, i.e., the additive quantization noise model (AQNM) [gersho2012vector, fletcher2007robust, orhan2015low] and the Bussgang decomposition [bussgang1952crosscorrelation]. Both approaches approximate the (nonlinear) quantization function using a linear model. However, in the literature, there are two distinct linear approximations referred to as the AQNM. The first is [gersho2012vector]
$Q(X)=X+q,$ (1)
where $Q(\cdot)$ and $q$ denote the quantization function and quantization error, respectively. The second is [fletcher2007robust]
$Q(X)=\alpha X+\eta,$ (2)
where $\alpha$ is a constant depending on the quantizers and on the distribution of $X$, and $\eta$ represents the quantization distortion (QD). Both (1) and (2) can be employed to analyze the worst-case system performance [diggavi2001worst, hassibi2003much] assuming that $q$ or $\eta$ is a Gaussian random variable uncorrelated with $X$. Model (2) was first derived in [fletcher2007robust] and named AQNM later in [orhan2015low]; it was also called the pseudo-quantization noise model in [zhang2016mixed]. Although (2) and the Bussgang decomposition were developed from separate technical lineages, it was shown in [demir2020bussgang] that the former is the latter tailored for the case of quantization. Therefore, we call the model in (2) as the Bussgang-based AQNM (BAQNM) while we refer to (1) as the AQNM for distinction. The AQNM is typically less accurate than the BAQNM because the assumption that $q$ is uncorrelated with $X$ is generally not satisfied. In contrast, $\eta$ is uncorrelated with $X$ based on the properties of the Bussgang decomposition. Furthermore, the QD covariance matrix is a key ingredient for the performance analysis and optimization with the BAQNM. A diagonal approximation of the QD covariance matrix was derived in [mezghani2012capacity, bai2013optimization], which has since then been widely used in the literature. Nonetheless, this approximation can introduce substantial error in the performance analysis and optimization in some scenarios [demir2020bussgang, prasad2020optimizing]. It has been numerically demonstrated that, at low SNR, using the BAQNM and the approximation of the QD covariance matrix can obtain a SE close to the channel capacity [orhan2015low, roth2017achievable]. However, at high SNR, the diagonal approximation of the QD covariance matrix results in non-negligible performance overestimation, especially for massive MIMO systems with very few ADC bits [roth2017achievable, demir2020bussgang].
I-B Contributions
Recently, Castañeda et. al [castaneda2021spawc, castaneda2021resolution] implemented the first fully digital systems with resolution-adaptive ADCs, where the resolution bits are dynamically adjusted to adapt to the instantaneous communications scenario, e.g., CSI and modulation scheme. They demonstrated that the adoption of resolution-adaptive ADCs can achieve power savings with several orders of magnitude, for realistic mmWave channels, underscoring the significant potential of this technology. Previous research on resolution-adaptive ADCs has mainly focused on bit allocation [bai2013optimization, choi2017resolution, nguyen2020energy] or the combined optimization of bit allocation with either transmit or receive beamforming design [ahmed2017joint, prasad2020optimizing, sheng2020energy]. However, the joint optimization of transmit-receive beamforming and bit allocation holds significant potential for achieving higher SE and providing valuable insights into the SE-EE tradeoff, yet it remains largely unexplored. Moreover, existing designs [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy] fail to adequately address the complex coupling between transmit-receive beamformers and the bit allocation vector, which demands a comprehensive and integrated design approach. This work bridges this critical gap by introducing a joint transmit-receive beamforming design and bit allocation in point-to-point MIMO systems employing resolution-adaptive ADCs. To facilitate the design, we analyze the BAQNM and the approximation of the QD covariance matrix. Table I outlines the distinctions between our work and prior studies. The specific contributions of this paper are summarized as follows.
  • We first present the fundamental properties of optimal quantization. Leveraging these properties and the Bussgang decomposition, we obtain a more accurate characterization of the BAQNM and the diagonal approximation of the QD covariance matrix, offering a new perspective compared to [mezghani2012capacity, bai2013optimization]. Our analysis shows that the BAQNM and the diagonal approximation of the QD covariance matrix only hold when Gaussian signals are optimally quantized. Furthermore, we reveal the connections between applying BAQNM and the arcsine law [jacovitti1994estimation] to one-bit quantization.
  • Building upon the theoretical findings explained above, we consider the joint transmit-receive beamforming design and bit allocation problem to maximize the SE subject to constraints on the transmit power budget and the total number of active ADC bits. This design problem is inherently complex due to its mixed-integer nature. We address it by first determining the beamformer under fixed ADC resolutions. Subsequently, we propose a low-complexity algorithm to iteratively optimize the ADC resolutions and the beamforming matrices.
  • We perform extensive numerical simulations based on the simulated QD covariance matrix, which has non-zero off-diagonal entries, using the proposed properties of optimal quantization. The results confirm the superiority of the proposed schemes over the state-of-the-art algorithms, especially beamforming alone with 22224444 ADC bits per RF chain. For example, in a $64\times 64$ MIMO system, the proposed design offers $6\%$ improvements in both SE and EE while requiring $40\%$ fewer active ADC bits compared with uniform bit allocation. Moreover, the SE-EE comparison shows that receiving more data streams with low-resolution ADCs can achieve higher SE and EE than receiving fewer data streams with high-resolution ADCs.
TABLE I: Comparison of prior works on resolution-adaptive ADCs.
Bit allocation
[bai2013optimization, choi2017resolution, nguyen2020energy]
Transmit beamforming [prasad2020optimizing]
Receive beamforming [ahmed2017joint, sheng2020energy]
Transmit-receive beamforming This work
I-C Organization and Notations
The rest of this paper is organized as follows. In Section II, we present the system and quantization models. The BAQNM and the approximation of the QD covariance matrix are then derived in Section III. We delve into the joint transmit-receive beamforming and bit allocation design in Section LABEL:sec:transceiver_design. Finally, we provide simulation results and conclusions in Sections LABEL:sec:simulation and LABEL:sec:conclusion, respectively.
Scalars, vectors, and matrices are denoted by the lowercase, boldface lowercase, and boldface uppercase letters, respectively. Furthermore, we use $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{T}}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{H}}$, and $\left(\cdot\right)^{-1}$ to represent the conjugate, transpose, conjugate transpose, and matrix inverse operators, respectively. $\left\|\cdot\right\|_{\mathcal{F}}$ signifies the Frobenius norm for matrices. In addition, the expectation and trace operators are represented by ${\mathbb{E}}\left(\cdot\right)$ and ${\rm Tr}\left(\cdot\right)$. We use $\left|a\right|$ and $\det\left(\mathbf{A}\right)$ to denote the absolute value of the scalar $a$ and the determinant of the matrix $\mathbf{A}$, respectively. The real and imaginary part operators are denoted by ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$, respectively. Moreover, ${\rm diag}(\mathbf{a})$ or ${\rm diag}(\mathbf{A})$ returns a diagonal matrix whose diagonal entries are the same as the elements of $\mathbf{a}$ or the diagonal entries of $\mathbf{A}$. In addition, $\mathbf{A}(:,1:J)$ represents the matrix consisting of the left $J$ columns of $\mathbf{A}$. Finally, we use $\mathbf{C}_{xy}$ and $\mathbf{C}_{x}$ to represent the cross-covariance matrix between $\mathbf{x}$ and $\mathbf{y}$ and the auto-covariance matrix of $\mathbf{x}$, respectively.
II System and Quantization Models
II-A System Model
We consider a point-to-point massive MIMO system where a transmitter (Tx) with $N_{\rm t}$ antennas communicates with a receiver (Rx) with $N_{\rm r}$ antennas. Here, “massive MIMO” refers to the deployment of large numbers of antennas at both the Tx and Rx [lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband]. We assume that the Tx is equipped with high-resolution digital-to-analog converters while resolution-adaptive ADCs are deployed at the Rx. The ADC bits at the Rx can be dynamically adjusted to adapt to the CSI. In practice, the resolution-adaptive ADCs can be fabricated by the flash architecture [yoo2002power, nahata2004high, rajashekar2008design]. Let $\mathbf{s}\in{\mathbb{C}}^{N_{\rm s}}$ ($N_{\rm s}\leq\min(N_{\rm t},N_{\rm r})$) be the transmitted signal vector of $N_{\rm s}$ data streams. We assume that $\mathbf{s}$ follows the Gaussian distribution and ${\mathbb{E}}[\mathbf{s}\mathbf{s}^{\scriptscriptstyle\mathsf{H}}]=\mathbf{I}$. Furthermore, let $\mathbf{F}\in{\mathbb{C}}^{N_{\rm t}\times N_{\rm s}}$ be the precoding matrix with the power constraint $\|\mathbf{F}\|_{\mathcal{F}}^{2}\leq P_{\rm t}$. Here, $P_{\rm t}$ denotes the transmit power budget of the Tx. The received signal (without quantization) at the Rx can be written as
$\mathbf{y}=\mathbf{H}\mathbf{F}\mathbf{s}+\mathbf{n},$ (3)
where $\mathbf{H}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ is the channel between the Tx and the Rx, and $\mathbf{n}$ denotes the additive white Gaussian noise (AWGN) vector, $\mathbf{n}\sim\mathcal{C}\mathcal{N}(0,\sigma_{\rm n}^{2}\mathbf{I})$, with $\sigma_{\rm n}^{2}$ being the noise power.
To characterize the system performance upper bound, we assume the availability of perfect CSI at both the Rx and Tx [mo2015capacity, ling2019performance, lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband]. We note that the channel can be assumed quasi-static in some point-to-point scenarios (e.g., wireless backhaul). In addition, channel estimation with adaptive-resolution ADCs was studied in [wang2022channel]. Furthermore, an ASIC receiver integrating both the resolution-adaptive ADCs and the channel estimation module was developed in [castaneda2021resolution]. The effect of imperfect CSI will be numerically investigated in Section LABEL:sec:simulation.
II-B Signal Model with Quantization
A scalar quantizer is fully characterized by its codebook $\mathcal{C}$ and threshold set $\mathcal{T}$. For a $b$-bit quantizer, we have $\mathcal{C}=\{c_{0},\ldots,c_{N_{\rm q}-1}\}$ and $\mathcal{T}=\{t_{0},\ldots,t_{N_{\rm q}}\}$, where $N_{\rm q}=2^{b}$ is the number of representation levels of the quantizer. Here, we assume $t_{0}=-\infty$ and $t_{N_{\rm q}}=\infty$, which allows inputs with arbitrary power. [^1] In practice, the input signal of ADCs outside the range $[t_{1},t_{N_{\rm q}-1}]$ can be clipped into the range of $[t_{1}-\iota,t_{N_{\rm q}-1}+\iota]$ where $\iota$ is an adjustable parameter depending on the constraints of hardware components, e.g., the automatic gain control (AGC). Let $Q(\cdot)$ denote the quantization function associated with $\mathcal{C}$ and $\mathcal{T}$. For a complex signal $x$, we have $Q(x)=Q(\Re\{x\})+jQ(\Im\{x\})$, with $Q(\Re\{x\})=c_{i},i\in\{0,\ldots,N_{\rm q}-1\}$ for $\Re\{x\}\in[t_{i},t_{i+1}]$. $Q(\Im\{x\})$ is obtained similarly.
The Bussgang decomposition can be applied to a vector space in the complex domain [demir2020bussgang]. Specifically, let $\mathbf{Q}:{\mathbb{C}}^{N}\rightarrow{\mathbb{C}}^{N}$ denote a scalar quantization function and $\mathbf{z}$ be the quantized output of $\mathbf{y}$ in (3). We can write $\mathbf{z}=\mathbf{Q}(\mathbf{y})$ or equivalently $z_{i}=Q_{i}(y_{i}),\;\forall i$, where $z_{i}$ and $y_{i}$ denote the $i$-th element of $\mathbf{z}$ and $\mathbf{y}$, respectively; $Q_{i}(\cdot)$ represents the associated quantization function. For the circularly-symmetric Gaussian random vector $\mathbf{y}$, the Bussgang decomposition implies
$\mathbf{z}=\mathbf{Q}(\mathbf{y})=\mathbf{G}\mathbf{y}+{\bm{\eta}},$ (4)
where $\mathbf{G}=\mathbf{C}_{zy}\mathbf{C}_{y}^{-1}$ denotes the Bussgang gain, and the distortion term ${\bm{\eta}}$ is uncorrelated to $\mathbf{y}$. In (4), ${\bm{\eta}}$ represents the QD vector with its covariance matrix given by
$\mathbf{C}_{\eta}={\mathbb{E}}[(\mathbf{z}-\mathbf{G}\mathbf{y})(\mathbf{z}-\mathbf{G}\mathbf{y})^{\scriptscriptstyle\mathsf{H}}]=\mathbf{C}_{z}-\mathbf{G}\mathbf{C}_{yz}.$ (5)
Furthermore, under some mild assumptions, the Bussgang gain $\mathbf{G}$ is shown to be diagonal, as detailed below.
Lemma 1 ([jacobsson2017quantized, bjornson2018hardware, demir2020bussgang])
Consider a circularly-symmetric Gaussian random vector $\mathbf{y}$ fed into scalar quantizers. With (4) modeling the quantization, we have $\mathbf{G}={\rm diag}\left(\mathbf{g}\right)$, where $g_{i}=\frac{{\mathbb{E}}[Q_{i}(y_{i})y_{i}^{*}]}{{\mathbb{E}}[|y_{i}|^{2}]}$ is the $i$-th element of $\mathbf{g}$.
Substituting (3) into (4), we obtain the quantized version of the signal received, expressed as
$\mathbf{z}=\mathbf{G}\mathbf{H}\mathbf{F}\mathbf{s}+\mathbf{e},$ (6)
where $\mathbf{e}=\mathbf{G}\mathbf{n}+{\bm{\eta}}$ represents the effective noise with covariance matrix $\mathbf{C}_{e}={\mathbb{E}}[\mathbf{e}\mathbf{e}^{\scriptscriptstyle\mathsf{H}}]=\mathbf{C}_{\eta}+\sigma_{\rm n}^{2}\mathbf{G}^{2}$. The post-combined signal at the Rx is expressed as
$\hat{\mathbf{s}}=\mathbf{U}^{\scriptscriptstyle\mathsf{H}}\mathbf{z}=\mathbf{U}^{\scriptscriptstyle\mathsf{H}}\mathbf{G}\mathbf{H}\mathbf{F}\mathbf{s}+\mathbf{U}^{\scriptscriptstyle\mathsf{H}}\mathbf{e},$ (7)
where $\mathbf{U}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ denotes the combining matrix. Although $\mathbf{s}$ is Gaussian distributed, $\mathbf{e}$ does not follow a Gaussian distribution because of the nonlinear quantization distortion. However, we can treat the effective noise vector $\mathbf{e}$ as a Gaussian random variable and obtain a lower bound of the SE as [hassibi2003much]
$R=\log\det\left(\mathbf{I}+(\mathbf{U}^{\scriptscriptstyle\mathsf{H}}\mathbf{C}_{e}\mathbf{U})^{-1}\mathbf{U}^{\scriptscriptstyle\mathsf{H}}\mathbf{G}\mathbf{H}\mathbf{F}\mathbf{F}^{\scriptscriptstyle\mathsf{H}}\mathbf{H}^{\scriptscriptstyle\mathsf{H}}\mathbf{G}\mathbf{U}\right).$ (8)
It is observed that the Bussgang gain $\mathbf{G}$ and the QD covariance matrix $\mathbf{C}_{\eta}$ are necessary for further analysis and optimization of the SE. For one-bit quantization, closed-form expressions for $\mathbf{G}$ and $\mathbf{C}_{\eta}$ can be derived based on the arcsine law. However, obtaining those for multi-bit quantization is significantly more challenging. A closed-form expression of $\mathbf{G}$ and a diagonal approximation of $\mathbf{C}_{\eta}$ were developed in [mezghani2012capacity, bai2013optimization] under the assumption that the quantizer satisfies the following properties:
${\mathbb{E}}[z_{i}-y_{i}]=0,$ (9)
${\mathbb{E}}[(z_{i}-y_{i})z_{i}]=0.$ (10)
However, the validity of these assumptions remains unclear, and thus the applicability of these results to general signal distributions and quantizers is uncertain. In the next section, we derive the BAQNM and diagonal approximation from a new perspective, aiming to clarify this uncertainty.
III Relations between Optimal Quantization and BAQNM
In this section, we first identify the fundamental properties of optimal quantizers. Then we leverage them to obtain the BAQNM and the approximation of the QD covariance matrix. Furthermore, we elaborate on the nuances between applying the BAQNM and the arcsine law to one-bit quantization.
III-A Properties of Optimal Quantizers
We first recall the definition of the optimal quantizer [max1960quantizing] below.
Definition 1 ([max1960quantizing])
Consider a real-valued random variable $X$. Let $f_{X}(x)$ denote its probability density function (PDF), and let $Q(x)=c_{i},i\in\{0,\ldots,N_{\rm q}-1\}$ be its quantized approximation for $x\in(t_{i},t_{i+1}]$. The mean square error (MSE) for the quantization can be expressed as
$D={\mathbb{E}}\left[\left(Q(x)-x\right)^{2}\right]=\sum\limits_{i=0}^{N_{\rm q}-1}\int_{t_{i}}^{t_{i+1}}(x-c_{i})^{2}f_{X}(x){\rm d}x.$ (11)
The optimal quantizer satisfies that its codebook and threshold set, i.e., $\{\mathcal{C},\mathcal{T}\}$, minimizes $D$.
Setting the derivatives of $D$ with respect to $t_{j}$ and $c_{j}$ to zeros yields
$t_{j}=\frac{c_{j}+c_{j-1}}{2},$ (12)
$c_{j}=\frac{\int_{t_{j}}^{t_{j+1}}xf_{X}(x){\rm d}x}{\int_{t_{j}}^{t_{j+1}}f_{X}(x){\rm d}x},$ (13)
which are referred to as the nearest neighbor condition and the centroid condition, respectively, [gersho2012vector, Ch. 6]. They are necessary for the optimal quantizer, also known as the Lloyd-Max quantizer [max1960quantizing] or the optimal non-uniform quantizer. The latter term reflects the fact that the representation levels of the optimal quantizer generally have non-uniform distribution in the real domain. In contrast, a uniform quantizer maintains equal distances between $c_{i}$ and $c_{i+1},\forall i$. With this constraint, the quantizer that minimizes $D$ in (11) is referred to as the optimal uniform quantizer.
Remark 1
The centroid condition requires that the representation level of each interval is its mean value. Mathematically, it can be written as [gersho2012vector]
${\mathbb{E}}[X|Q(X)]=Q(X),$ (14)
which was used in [fletcher2007robust] as a basic assumption for deriving the model (2). Therefore, the BAQNM is limited to the optimal quantizer.
For a specific input signal, we can employ the Lloyd-Max algorithm [max1960quantizing] that iteratively updates $\mathcal{T}$ and $\mathcal{C}$ based on (12) and (13) to find the optimal quantizer. However, this iterative algorithm results in a high time complexity, especially for high-resolution quantization. We herein present an efficient approach to obtain the optimal quantization for Gaussian signals with the proposition below.
Proposition 1
Let $X$ be a real-valued, zero-mean, and unit-variance random variable, and let $Y=\sigma_{\rm y}X$. Then, we have
$Q_{\rm y}(Y)=\sigma_{\rm y}Q_{\rm x}(X)=\sigma_{\rm y}Q_{\rm x}\left(\frac{Y}{\sigma_{\rm y}}\right),$ (15)
$\gamma={\mathbb{E}}\left[(Q_{\rm x}(X)-X)^{2}\right]=\frac{{\mathbb{E}}\left[(Q_{\rm y}(Y)-Y)^{2}\right]}{\sigma_{\rm y}^{2}},$ (16)
where $Q_{\rm y}(Y)$ and $Q_{\rm x}(X)$ denote the optimal quantized output of $Y$ and $X$, respectively.
Proof:
See Appendix LABEL:prof:scaling_and_distortion_invariance. ∎
We refer to $\gamma$ as the distortion factor and the properties in (15) and (16) as the scaling law and distortion invariance, respectively. The scaling law enables a convenient way to obtain the optimal quantization for any Gaussian signal with a known variance. For example, we can obtain the optimal element-wise quantization of the received signal vector $\mathbf{y}$ in (3) with covariance matrix
$\mathbf{C}_{y}={\mathbb{E}}[\mathbf{y}\mathbf{y}^{\scriptscriptstyle\mathsf{H}}]=\mathbf{H}\mathbf{F}\mathbf{F}^{\scriptscriptstyle\mathsf{H}}\mathbf{H}^{\scriptscriptstyle\mathsf{H}}+\sigma_{\rm n}^{2}\mathbf{I}$ (17)
based on the optimal quantizer for the standard Gaussian signal [max1960quantizing]. Regarding the distortion factor, we note the following property.
Proposition 2
For a zero-mean complex random variable $X=\Re\{X\}+j\Im\{X\}$ with variance $\sigma_{X}^{2}$, assume that $\Re\{X\}$ and $\Im\{X\}$ are independent and identically distributed (i.i.d.) with the same variance $\frac{\sigma_{X}^{2}}{2}$ and are independently quantized by two identical Lloyd-Max quantizers $Q(\cdot)$. With $\chi=Q(X)-X$, we obtain
${\mathbb{E}}[Q(X)]={\mathbb{E}}[X],$ (18)
${\mathbb{E}}[Q(X)\chi^{*}]=0,$ (19)
$\gamma=\frac{{\mathbb{E}}[\left|\chi\right|^{2}]}{{\mathbb{E}}[\left|X\right|^{2}]}=\frac{{\mathbb{E}}[\Re\{\chi\}^{2}]}{{\mathbb{E}}[\Re\{X\}^{2}]}=\frac{{\mathbb{E}}[\Im\{\chi\}^{2}]}{{\mathbb{E}}[\Im\{X\}^{2}]}.$ (20)

### v3

@@
 Furthermore, mixed-ADC systems, which simultaneously deploy one-bit and high-resolution ADCs, have been shown to perform better than fixed-resolution architectures, especially at high SNR [zhang2016mixed, zhang2017performance, pirzadeh2018spectral].
-On the other hand, resolution-adaptive ADCs have been studied in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, prasad2020optimizing, castaneda2021resolution, castaneda2021spawc, sheng2020energy] to flexibly balance the SE-EE tradeoff of low-resolution systems.
-For instance, it has been shown in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, sheng2020energy] that efficient bit allocation strategies can offer a higher EE compared with uniform-resolution architectures.
+On the other hand, resolution-adaptive ADCs have been studied in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, prasad2020optimizing, castaneda2021resolution, castaneda2021spawc, sheng2020energy, verenzuela2021optimal, verenzuela2017per] to flexibly balance the SE-EE tradeoff of low-resolution systems.
+For instance, it has been shown in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, sheng2020energy, verenzuela2021optimal, verenzuela2017per] that efficient bit allocation strategies can offer a higher EE compared with uniform-resolution architectures.
 Moreover, Castañeda et al.
@@
 where $\alpha$ is a constant depending on the quantizers and on the distribution of $X$, and $\eta$ represents the quantization distortion (QD).
-Both (1) and (2) can be employed to analyze the worst-case system performance [diggavi2001worst, hassibi2003much] assuming that $q$ or $\eta$ is a Gaussian random variable uncorrelated with $X$.
+Both (1) and (2) can be employed to analyze the worst-case system’s performance [diggavi2001worst, hassibi2003much] assuming that $q$ or $\eta$ is a Gaussian random variable uncorrelated with $X$.
 Model (2) was first derived in [fletcher2007robust] and named AQNM later in [orhan2015low]; it was also called the pseudo-quantization noise model in [zhang2016mixed].
@@
 al [castaneda2021spawc, castaneda2021resolution] implemented the first fully digital systems with resolution-adaptive ADCs, where the resolution bits are dynamically adjusted to adapt to the instantaneous communications scenario, e.g., CSI and modulation scheme.
-They demonstrated that the adoption of resolution-adaptive ADCs can achieve power savings with several orders of magnitude, for realistic mmWave channels, underscoring the significant potential of this technology.
-Previous research on resolution-adaptive ADCs has mainly focused on bit allocation [bai2013optimization, choi2017resolution, nguyen2020energy] or the combined optimization of bit allocation with either transmit or receive beamforming design [ahmed2017joint, prasad2020optimizing, sheng2020energy].
+They demonstrated that the adoption of resolution-adaptive ADCs can achieve power savings with several orders of magnitude for realistic mmWave channels, underscoring the significant potential of this technology.
+Previous research on resolution-adaptive ADCs has mainly focused on bit allocation [bai2013optimization, choi2017resolution, nguyen2020energy, verenzuela2021optimal, verenzuela2017per] or the combined optimization of bit allocation with either transmit or receive beamforming design [ahmed2017joint, prasad2020optimizing, sheng2020energy].
 However, the joint optimization of transmit-receive beamforming and bit allocation holds significant potential for achieving higher SE and providing valuable insights into the SE-EE tradeoff, yet it remains largely unexplored.
-Moreover, existing designs [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy] fail to adequately address the complex coupling between transmit-receive beamformers and the bit allocation vector, which demands a comprehensive and integrated design approach.
+Moreover, existing designs [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy, verenzuela2021optimal, verenzuela2017per] fail to adequately address the complex coupling between transmit-receive beamformers and the bit allocation vector, which demands a comprehensive and integrated design approach.
 This work bridges this critical gap by introducing a joint transmit-receive beamforming design and bit allocation in point-to-point MIMO systems employing resolution-adaptive ADCs.
-To facilitate the design, we analyze the BAQNM and the approximation of the QD covariance matrix.
+To facilitate the design, it is essential to properly model the quantization process.
+Although the BAQNM is widely adopted in the literature, its foundational assumptions, which significantly influence its accuracy, have not been thoroughly examined.
+Existing studies mainly use the BAQNM as a plug-and-play tool without evaluating its applicability, which may lead to questionable conclusions.
+To address this issue, we aim to develop a more accurate quantization model that enables reliable and insightful analysis.
 Table I outlines the distinctions between our work and prior studies.
 The specific contributions of this paper are summarized as follows.
-  • We first present the fundamental properties of optimal quantization.
-Leveraging these properties and the Bussgang decomposition, we obtain a more accurate characterization of the BAQNM and the diagonal approximation of the QD covariance matrix, offering a new perspective compared to [mezghani2012capacity, bai2013optimization].
+TABLE I: Comparison of prior works on resolution-adaptive ADCs.
+Reference Bit allocation Transmit beamforming Receive beamforming Quantization modeling analysis
+[bai2013optimization, choi2017resolution, nguyen2020energy, verenzuela2021optimal, verenzuela2017per]
+[prasad2020optimizing]
+[ahmed2017joint, sheng2020energy]
+This work
+  • We first present the fundamental properties of optimal quantization, including the scaling law, distortion invariance, and some important statistical properties between random variable and its quantized output.
+The scaling law enables efficient derivation of the optimally quantized outputs for Gaussian signals with different variances, significantly reducing computational and time complexities.
+As a result, it allows to efficiently obtain the actual QD covariance matrix via numerical methods, which includes the non-zero off-diagonal entries.
+This leads to more reliable performance evaluation compared to that solely relying on the theoretical diagonal approximation.
+In addition, the distortion invariance property ensures that the Bussgang gain matrix in the BAQNM depends solely on the distortion factor, which is determined by the quantizer resolution.
+We propose a more accurate expression of this distortion factor, enabling reliable performance evaluation when employing the BAQNM.
+  • Leveraging the fundamental properties of optimal quantization and the Bussgang decomposition, we obtain a more accurate characterization of the BAQNM and the diagonal approximation of the QD covariance matrix compared to those developed in [mezghani2012capacity, bai2013optimization].
 Our analysis shows that the BAQNM and the diagonal approximation of the QD covariance matrix only hold when Gaussian signals are optimally quantized.
-Furthermore, we reveal the connections between applying BAQNM and the arcsine law [jacovitti1994estimation] to one-bit quantization.
+Moreover, we reveal the connections and nuances between applying BAQNM and the arcsine law [jacovitti1994estimation] to one-bit quantization.
+The BAQNM applies to only the optimal quantization, whereas the arcsine law implies more general quantization for one-bit systems.
   • Building upon the theoretical findings explained above, we consider the joint transmit-receive beamforming design and bit allocation problem to maximize the SE subject to constraints on the transmit power budget and the total number of active ADC bits.
@@
 Subsequently, we propose a low-complexity algorithm to iteratively optimize the ADC resolutions and the beamforming matrices.
-  • We perform extensive numerical simulations based on the simulated QD covariance matrix, which has non-zero off-diagonal entries, using the proposed properties of optimal quantization.
+  • We perform extensive numerical simulations based on the simulated QD covariance matrix, which has non-zero off-diagonal entries, using the scaling law of optimal quantization.
 The results confirm the superiority of the proposed schemes over the state-of-the-art algorithms, especially beamforming alone with 22224444 ADC bits per RF chain.
@@
 Moreover, the SE-EE comparison shows that receiving more data streams with low-resolution ADCs can achieve higher SE and EE than receiving fewer data streams with high-resolution ADCs.
-TABLE I: Comparison of prior works on resolution-adaptive ADCs.
-Bit allocation
-[bai2013optimization, choi2017resolution, nguyen2020energy]
-Transmit beamforming [prasad2020optimizing]
-Receive beamforming [ahmed2017joint, sheng2020energy]
-Transmit-receive beamforming This work
 I-C Organization and Notations
@@
 Furthermore, we use $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{T}}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{H}}$, and $\left(\cdot\right)^{-1}$ to represent the conjugate, transpose, conjugate transpose, and matrix inverse operators, respectively.
-$\left\|\cdot\right\|_{\mathcal{F}}$ signifies the Frobenius norm for matrices.
+$\left\|\cdot\right\|_{\mathcal{F}}$ signifies the Frobenius norm for matrices, whereas $\otimes$ denotes the Kronecker product.
 In addition, the expectation and trace operators are represented by ${\mathbb{E}}\left(\cdot\right)$ and ${\rm Tr}\left(\cdot\right)$.
@@
 In addition, $\mathbf{A}(:,1:J)$ represents the matrix consisting of the left $J$ columns of $\mathbf{A}$.
-Finally, we use $\mathbf{C}_{xy}$ and $\mathbf{C}_{x}$ to represent the cross-covariance matrix between $\mathbf{x}$ and $\mathbf{y}$ and the auto-covariance matrix of $\mathbf{x}$, respectively.
+Lastly, we use $\mathbf{C}_{xy}$ and $\mathbf{C}_{x}$ to represent the cross-covariance matrix between $\mathbf{x}$ and $\mathbf{y}$ and the auto-covariance matrix of $\mathbf{x}$, respectively.
 II System and Quantization Models
@@
 where $\mathbf{H}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ is the channel between the Tx and the Rx, and $\mathbf{n}$ denotes the additive white Gaussian noise (AWGN) vector, $\mathbf{n}\sim\mathcal{C}\mathcal{N}(0,\sigma_{\rm n}^{2}\mathbf{I})$, with $\sigma_{\rm n}^{2}$ being the noise power.
-To characterize the system performance upper bound, we assume the availability of perfect CSI at both the Rx and Tx [mo2015capacity, ling2019performance, lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband].
+To characterize the upper bound on the system’s performance, we assume the availability of perfect CSI at both the Rx and Tx [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy, mo2015capacity, ling2019performance, lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband].
+The effect of imperfect CSI will be numerically investigated in Section LABEL:sec:imperfect_implementation.
 We note that the channel can be assumed quasi-static in some point-to-point scenarios (e.g., wireless backhaul).
 In addition, channel estimation with adaptive-resolution ADCs was studied in [wang2022channel].
-Furthermore, an ASIC receiver integrating both the resolution-adaptive ADCs and the channel estimation module was developed in [castaneda2021resolution].
-The effect of imperfect CSI will be numerically investigated in Section LABEL:sec:simulation.
-II-B Signal Model with Quantization
+Furthermore, an ASIC receiver integrating both the resolution-adaptive ADCs and the channel estimation module was implemented in [castaneda2021resolution].
+II-B Channel Model
+The mmWave propagation environment can be well characterized by the Saleh-Valenzuel channel model [rangan2014millimeter].
+Assuming a uniform planar array (UPA) at both the Tx and the Rx, the channel matrix is expressed as [yu2016alternating, lin2019transceiver, qi2022hybrid, zhang2017hybridly]
+$\mathbf{H}=\sqrt{\frac{N_{\rm t}N_{\rm r}}{N_{\rm cl}N_{\rm ray}}}\sum\limits_{i=1}^{N_{\rm cl}}\sum\limits_{l=1}^{N_{\rm ray}}\alpha_{il}\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)\mathbf{a}_{\rm t}^{\scriptscriptstyle\mathsf{H}}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right),$ (4)
+where $N_{\rm cl}$ and $N_{\rm ray}$ indicate the number of clusters and distinct rays within each cluster, respectively, and $\alpha_{il}$ denotes the gain of $l$-th ray in the $i$-th propagation cluster.
+In addition, $\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)$ and $\mathbf{a}_{\rm t}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right)$ represent the receive and transmit array response vectors, respectively, where $\theta_{il}^{r}$ ($\phi_{il}^{\rm r}$) and $\theta_{il}^{\scriptscriptstyle\mathsf{T}}$ ($\phi_{il}^{\rm t}$) stand for the azimuth (elevation) angles of arrival/departure (AoAs/AoDs) of the $l$-th ray in the $i$-th propagation cluster, respectively.
+Assume that the Rx deploys a UPA of size $N_{\rm r,h}\times N_{\rm r,v}$ with $N_{\rm r}=N_{\rm r,h}N_{\rm r,v}$.
+Defining $\rho_{il}^{\rm r}=\sin(\theta_{il}^{\rm r})\sin(\phi_{il}^{\rm r})$ and $\varrho_{il}^{\rm r}=\cos(\theta_{il}^{\rm r})$, the array response vector $\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)$ at the Rx can be expressed as[yu2016alternating, lin2019transceiver]
+$\mathbf{a}_{\rm r}\left(\rho_{il}^{\rm r},\varrho_{il}^{\rm r}\right)=\mathbf{a}_{\rm r,h}\left(\rho_{il}^{\rm r}\right)\otimes\mathbf{a}_{\rm r,v}\left(\varrho_{il}^{\rm r}\right),$ (5)
+with
+$\mathbf{a}_{\rm r,h}\left(\rho_{il}^{\rm r}\right)=\frac{1}{\sqrt{N_{\rm r,h}}}[1,e^{j\pi\rho_{il}^{\rm r}},\cdots,e^{j(N_{\rm r,h}-1)\pi\rho_{il}^{\rm r}}]^{\scriptscriptstyle\mathsf{T}},$ (6)
+$\mathbf{a}_{\rm r,v}\left(\varrho_{il}^{\rm r}\right)=\frac{1}{\sqrt{N_{\rm r,v}}}[1,e^{j\pi\varrho_{il}^{\rm r}},\cdots,e^{j(N_{\rm r,v}-1)\pi\varrho_{il}^{\rm r}}]^{\scriptscriptstyle\mathsf{T}}.$
+The array response vector $\mathbf{a}_{\rm t}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right)$ at the Tx can be modeled similarly.
+II-C Signal Model with Quantization
 A scalar quantizer is fully characterized by its codebook $\mathcal{C}$ and threshold set $\mathcal{T}$.
@@
 where $\mathbf{U}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ denotes the combining matrix.
-Although $\mathbf{s}$ is Gaussian distributed, $\mathbf{e}$ does not follow a Gaussian distribution because of the nonlinear quantization distortion.
+Although $\mathbf{s}$ is Gaussian distributed, $\mathbf{e}$ does not follow a Gaussian distribution because of the nonlinear QD.
 However, we can treat the effective noise vector $\mathbf{e}$ as a Gaussian random variable and obtain a lower bound of the SE as [hassibi2003much]
@@
 Proposition 1
-Let $X$ be a real-valued, zero-mean, and unit-variance random variable, and let $Y=\sigma_{\rm y}X$. Then, we have
+Let $X$ be a real-valued, zero-mean, and unit-variance Gaussian random variable, and let $Y=\sigma_{\rm y}X$. Then, we have
 $Q_{\rm y}(Y)=\sigma_{\rm y}Q_{\rm x}(X)=\sigma_{\rm y}Q_{\rm x}\left(\frac{Y}{\sigma_{\rm y}}\right),$ (18)
@@
 Proposition 2
-For a zero-mean complex random variable $X=\Re\{X\}+j\Im\{X\}$ with variance $\sigma_{X}^{2}$, assume that $\Re\{X\}$ and $\Im\{X\}$ are independent and identically distributed (i.i.d.) with the same variance $\frac{\sigma_{X}^{2}}{2}$ and are independently quantized by two identical Lloyd-Max quantizers $Q(\cdot)$.
+For a zero-mean complex random variable $X=\Re\{X\}+j\Im\{X\}$ with variance $\sigma_{X}^{2}$, assume that $\Re\{X\}$ and $\Im\{X\}$ are independent and identically distributed (i.i.d.) with the same variance $\frac{\sigma_{X}^{2}}{2}$ and are independently quantized with two identical Lloyd-Max quantizers $Q(\cdot)$.
 With $\chi=Q(X)-X$, we obtain